Neutron stars consist only of neutrons and have unbelievably high densities. A typical mass and radius for a neutron star might be 2.65E+28 kg and 1.12E+3 m. Calculate the density of such a star.
Answer #1

Given, m= 2.65 x 1025 kg ; R = 1-12 x 10 m. weknow, Density (L) = mass (m) Volume (V) f = m ITR3 3 > f = 2.6581028 3 m 3 xr x